Circular No. 7196

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

GRS 1915+105 E. Gérard, Observatoire de Paris-Meudon, communicates: “The large

flare reported on June 8 (IAUC 7195) is still in progress on June 9 at 1410 and 3300 MHz. The flux densities measured with the Nançay radio telescope are respectively 0.2 and 0.4 Jy (June 8.1 UT) and 0.5 and 0.5 Jy (June 9.1). In addition, the 3300-MHz flux density varied by nearly a factor of 2 during 40 min on June 9.1, suggesting that quasiperiodic oscillations are superposed on the slower flare variation.”

V382 VELORUM M. Orio, Turin Observatory and University of Wisconsin; V. Torroni,

BeppoSAX Science Operation Center, Rome; and R. Ricci, BeppoSAX Science Data Center, Rome, report on behalf of the BeppoSAX team: “BeppoSAX observed N Vel 1999 from June 7.622 to 8.538 UT. Preliminary analysis of the data show that the source is relatively strong in x-rays with a count rate of 0.15 counts/s in the two MECS instruments (1.6–10 keV) and 0.05 counts/s in the LECS instrument (0.1–10 keV), corresponding to a flux at 2–10 keV of 10^{−11} erg cm^{−2} s^{−1}. At a distance of ∼ 2000 pc (Della Valle and Shore 1999, private communication), this implies a luminosity ∼ 1.9 × 10^{33} erg/s. The x-ray emission seems to be thermal and due to hot shocked gas, as in the case of the x-ray detection of V838 Her, five days after the outburst (Lloyd et al. 1991, Nature 356, 222). From the LECS/MECS ratio, we find that the presence of a supersoft component should be relatively small, if any. A more detailed spectral analysis will be performed in the next few days.”

Additional photometry of N Vel 1999 by A. C. Gilmore (cf. IAUC 7179) at Mount John (reference star HD 92518, with assumed V = 6.88, U − B = +0.03, B − V = +0.02, V − R = +0.02, V − I −0.02): June 6.484 UT, V = 6.06, U − B = −0.96, B − V = −0.03, V − R = +1.20, V − I = +0.90.

SUPERNOVAE 1999cd AND 1999cf R. Szabo reports CCD R magnitude estimates of two supernovae using

the 1-m telescope at Piszkéstető, Hungary (based on USNO-A1.0 R magnitudes): SN 1999cd in NGC 3646, May 25.944 UT, 17.0; SN 1999cf in UGC 8539, May 25.837, 16.0.

1999 June 9  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green